OpenStudy (anonymous):

x^2/x+3 -9/x+3 they have the same denominators but how do I subtract the top

OpenStudy (anonymous):

\[\frac{x^2}{x+3}-\frac{9}{x+3}=\frac{x^2-9}{x+3}\] however you are not done, as you can factor the numerator and then cancel the common factor

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[x^2-9\] is the difference of two squares, so it factors as \[x^2-9=(x-3)(x+3)\]

OpenStudy (anonymous):

so x-3

OpenStudy (anonymous):

final answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome thanks

OpenStudy (anonymous):

or if you want to impress your math teacher, you can write \(x-3\) if \(x\neq- 3\)

OpenStudy (anonymous):

yw